The average weight of pine trees in a forest is normally distributed with mean 53.45 and

a standard deviation of 2.51. A sample of 20 pine trees are taken from the forest and their mean

is recorded 64.85. Test the hypothesis that the average weight is less than 2.51. Using Z 2.83  = .


1
Expert's answer
2021-02-25T00:47:06-0500

Given, mean, "\mu=53.45"

standard deviation ,"\sigma=2.51"

Sample size, "n=20"

Mean of trees, "x=64.85"


The z-value is given by-


"z=\dfrac{x-\mu}\sigma"


"=\dfrac{64.85-53.45}{2.51}"



"=\dfrac{11.4}{2.51}=4.54"


Since The value of z Calculated is more than 2.51.


So Hypothesis is failed.